OpenStudy (anonymous):

Which property of equality can be used to solve the equation n + 9 = -3? Addition Subtraction Multiplication Division

OpenStudy (anonymous):

well how would you solve this equation?

OpenStudy (anonymous):

or what needs to be done for n to be by itself?

OpenStudy (anonymous):

Subtraction

OpenStudy (uri):

n + 9 = -3 n=-3-9 n=-12 Now what do you think which property this may be?

OpenStudy (anonymous):

It seems to me that it is subtraction

OpenStudy (uri):

:)